OpenStudy (anonymous):

Lisa has 5 more green marbles than blue marbles. She has a total of 40 green and blue marbles. Which system of equations represents this situation if x is the number of green marbles and y is the number of blue marbles? A: {y = x+5 {x+y=40 B: {x=y+5 {x+y=40 C: {y=x+5 {y=x+40 D: {x = y + 5 {x=y+40

OpenStudy (anonymous):

.............she should buy some red marbles.

OpenStudy (anonymous):

green marbles = blue marbles + 5. green marbles + blue marbles = 40. substitute x and y. for 'green marbles' and 'blue marbles'

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

How would it be b? d makes alot more sense

OpenStudy (anonymous):

given: x=green marbles y=blue marbles x=y+5 (since, green marbles is great than blue marbles by 5) x+y=40 (since the total number of marbles is 40) equation: x=y+5 x+y=40

OpenStudy (anonymous):

^ agree. mertsj, show working.

OpenStudy (anonymous):

"a total of 40 green and blue marbles" makes D wrong.

OpenStudy (anonymous):

tnx for the medal. .

OpenStudy (anonymous):

B.